OpenStudy (anonymous):

James has 35 coins in his pocket, all of them dimes and quarters. If he has a total of $5.15, how many quarters does he have?

OpenStudy (pfenn1):

If x = number of quarters, then number of dimes = 35-x. And we also know that The number of dimes times $0.10 plus the number of quarters times $0.25 equals $5.15.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so do i do 5.15/.25??

OpenStudy (binary3i):

sorry connection problem don't block me.

OpenStudy (pfenn1):

*number

OpenStudy (anonymous):

.10d+.25q

OpenStudy (pfenn1):

Close, you just need to finish that equation. Your sum adds up to 5.15. 0.10d + 0.25q = 5.15 We also know that the number of dimes and the number of quarters add up to 35. d +q = 35

OpenStudy (anonymous):

yeah

OpenStudy (pfenn1):

Do you understand what I did?

OpenStudy (anonymous):

nope

OpenStudy (pfenn1):

We know that James has 35 coins in his pocket. Some are dimes and some are quarters. So the number of dimes plus the number of quarters must equal 35. d + q = 35 Do you understand how I derived that equation?

OpenStudy (anonymous):

i figured it out there are 11 quarter thank you :)

OpenStudy (pfenn1):

Great. That is the answer I got as well.